OpenStudy (anonymous):

(dy/dx)+(y/3)=0 How do I get the y variables on one side? Thanks!

OpenStudy (unicornpoopcookies):

You subtract them from one side to move it :)

OpenStudy (abb0t):

\(\frac{ dy }{ dx } = \frac{ y }{ 3 } \rightarrow \frac{ dy }{ y } = \frac{ 1 }{ 3 }dx\) now integrate

OpenStudy (anonymous):

so 3dy + ydx to start?

OpenStudy (abb0t):

it should be negative \(\frac{1}{3}\) by th way

OpenStudy (anonymous):

How can the y/3 turn into 1/3 I don't get it

OpenStudy (anonymous):

well -1/3

OpenStudy (abb0t):

you wan to get dy annd 'y' together, and dx and'x' toether so u can integrate

OpenStudy (abb0t):

all I did was play around with it using algebra.

OpenStudy (anonymous):

I am so bad at algebra, I see what you did there, thank you sooo much!

OpenStudy (abb0t):

@Luis_Rivera #swag

OpenStudy (anonymous):

Never take a 15 year break from mathmatics

OpenStudy (anonymous):

cause you will spell even the word wrong! lol